MIT-REPORT

Synchronization and Pipeline Design for a Multithreaded Massively Parallel

Computer

Shuichi Sakai

Computation Structure Group MIT-LCS

Acknowledgment

Dr. Sakai would like to thank

• Prof. Arvind, Prof. Papadopoulos, and Dr. G. A. Boughton for supporting his research (also his life),

• the members of CSG for the fruitful discussions (also for buying his items and carrying his boxes)

• Prof. Dennis and Dr. Nikhil for his helpful comments

• Ms. Maderer for her support.

Introduction

This presents the optimization of a synchronization and a pipeline in a multithreaded massively parallel computer.

Strategies

• study the past and the present

• utilize, improve or forget

• reconsider ourselves Remark: Here, we concentrate on data-driven sync. and a processor pipeline based on a dynamic model.

Proposed Data-Driven Synchronizations

• Hashing in Manchester Machine and SIGMA-1

• ETS in the Monsoon

• Direct Matching in the EM-4

• Direct Match in Sandia Dataflow

• *T Synchronization

• Advanced Direct Matching in the EM-5

Matching by Pseudo Associative Mechanisms

Advantages

• high spatial efficiency

• flexibility Drawbacks

• complex

• pipeline bubbles caused by a hash-miss • difficult to understand what is happening